OpenStudy (anonymous):

1/3 is how many 12ths

OpenStudy (bahrom7893):

\[\frac{1}{3}=\frac{1*4}{3*4}=\frac{4}{12}=4*\frac{1}{12}\]

OpenStudy (bahrom7893):

Thus there are 4 12ths there